OpenStudy (anonymous):

Derivative x/(√1-x^2)

OpenStudy (abb0t):

Is the function: \(\large \frac{x}{\sqrt{1-x^2}}\) OR is it: \(\large \frac{x}{\sqrt{1}-x^2}\)

OpenStudy (anonymous):

The 1st one

OpenStudy (anonymous):

quotient rule for this one \[\left(\frac{f}{g}\right)'=\frac{f'g-g'f}{g^2}\] with \[f(x)=x,f'(x)=1,g(x)=\sqrt{1-x^2}, g'(x)=\frac{-x}{\sqrt{1-x^2}}\]

OpenStudy (abb0t):

Why are you introducing integration and substitution if he is struggling with derivatives?! @SandeepReddy

OpenStudy (anonymous):

@abb0t apologies i will remove it.